# Compression of quantum shallow-circuit states

### v1

@@ I.3 Discussions
This work serves as the first step of establishing a new direction of coherent quantum information processing where the complexity of resources determines the rate and performance of processing, which goes beyond the existing literature that focused on incoherent information processing [9, 10, 11]. For future perspectives, it is our goal to consider more tasks such as cloning [19, 20, 21, 22] and gate programming [35, 36, 37, 38, 39] and, ultimately, to re-examine the entire quantum Shannon theory established in the past decade from the new perspective of the NISQ era.
The remaining part of this paper is arranged as follows. In Section II, we prepare necessary preliminaries for later sections. In Section III, we introduce a new task of efficient local-parameterization of quantum states, which will be the basis of our compression protocol. In Section IV, we develop a technique of local asymptotic equivalence between $N$-copy shallow-circuit states and multi-mode coherent states of light, which will be a key subroutine of the compression protocol. Finally, in Section V, we design the compression protocol for shallow-circuit states, evaluate its performance, and show its key features.
II Preliminaries
II.1 Conventions and notations
@@
II.2 Distance measures for quantum states and gates
@@
where $\mathcal{H}_{A}\simeq\mathcal{H}$ is an ancillary space. For two unitary channels $\mathcal{U}$ and $\mathcal{V}$, the diamond norm is well-captured by the operator norm (cf. [40, Proposition 1.6]):
@@
II.3 Properties of $N$-copy states.
@@
Therefore, the task of $N$-copy state compression cannot be trivially reduced to quantum state tomography. In Subsection V.6, we will show the same result for shallow-circuit states.
III Efficient local-parameterization of shallow-circuit states
In Section V we will see that, to achieve faithful compression, we need to pin the shallow-circuit state $|\psi_{{\rm sc}}\rangle$ down to a small local region parameterised by up to $\mathsf{poly}(n)$ real parameters. We name such a task as local-parameterization of quantum states, which is achieved by measuring copies of $|\psi_{{\rm sc}}\rangle$ and then constructing the local model. However, it is more challenging than the task of quantum state tomography, as one has to not only identify a neighborhood of the state but also parameterise this neighborhood efficiently.
III.1 Local-parameterization of quantum states
@@
Lemma 2 (Local-parameterization of $\epsilon$-rotations).
@@
Definition 2 (Local-parameterization of quantum states).
Given a set $\mathsf{S}=\{|\psi\rangle\}$ of $n$-qubit quantum states and a fixed quantum state $|\psi_{0}\rangle=U_{0}|0\rangle^{\otimes n}$ with the guarantee that $d_{\operatorname{Tr}}(\psi,\psi_{0})\leq\epsilon$ for every $|\psi\rangle\in\mathsf{S}$, the task of local-parameterization of $\mathsf{S}$ is to identify a set $\mathsf{W}_{\rm loc}$ of $O(\epsilon)$-rotations, such that for every $|\psi\rangle\in\mathsf{S}$ there exists a $W\in\mathsf{W}_{\rm loc}$ with $|\psi\rangle=U_{0}W|0\rangle^{\otimes n}$. Moreover, the local-parameterization is efficient if $\mathsf{W}_{\rm loc}$ can be parameterized with $\mathsf{poly}(n)$ local parameters, i.e., if there exists an on-to function $\mathcal{P}:E^{f}\to\mathsf{W}_{\rm loc}$ with $E\subset\mathbb{R}$ being an $O(\epsilon)-$ length interval and $f\in\mathsf{poly}(n)$.
@@
The above “vanilla” local-parameterization of quantum states appears a simple task. However, making it parameter-efficient is not a trivial task. The local-parameterization in Lemma 3 consists of generic $n$-qubit unitaries and has exponentially many degrees of freedom.
In practice, it is more favorable to seek an efficient localization. A hope is when the set of states under consideration has certain structure. For instance, when $\mathsf{S}$ consists of states generated by acting a low-dimensional unknown quantum gate upon a fixed state. If all such states in $\mathsf{S}$ are close to the original fixed state, it is quite natural to speculate that we could probably replace this unknown unitary by an $O(\epsilon)$-rotation to achieve an efficient local-parameterization (since the $O(\epsilon)$-rotation is also low-dimensional). Somewhat surprisingly, this intuition turns out to be wrong:
@@
In general, it is not possible to partially local-parameterize a quantum state. That is, there exists a bipartite pure state $|\psi\rangle_{AB}$ and a unitary $\mathcal{V}_{B}$ acting on $B$ such that:
@@
The purpose of the above remark is to justify that our goal – an efficient local-parameterization of shallow-circuit states – is not as simple as it appears. One of our main results is to rigorously prove that such an efficient local-parameterization indeed exists for shallow-circuit states.
Theorem 1 (Efficient local-parameterization of shallow-circuit states.).
@@
Here $W=\prod_{j=1}^{n_{\rm gate}}W^{28\epsilon}_{j}$ with $n_{\rm gate}\leq n/d+4$, and each $W^{28\epsilon}_{j}$ is a $(28\epsilon)$-rotation acting on no more than $8d$ qubits. Among $\{W_{j}^{28\epsilon}\}$, the maximum of the number of gates that act nontrivially on the same qubit is upper bounded by $n_{\rm overlap}=4$. The set $\mathsf{W}_{\rm loc}$ can be parameterized with $O(n\cdot(2^{16d}/d))$ (real) local parameters.
Note that the bound on the number of parameters can be made tighter (but the bound might become less succinct). Details can found in the proof; see Subsection III.2.
@@ III.2 Proof of Theorem 1
Grouping qubits. For convenience of discussion, we group the $n$ qubits into subsets and label them using the following convention. We define a function $Q=Q(U)$ that maps a unitary $U$ to the qubits that it acts nontrivially on. In particular, $Q_{j}^{(i)}:=Q(U_{j}^{(i)})$ is the collection of qubits that $U_{j}^{(i)}$ acts non-trivially on. We further divide $Q_{j}^{(i)}$ into subsets $Q_{j}^{(i)}(T)$ (top), $Q_{j}^{(i)}(M)$ (middle) and $Q_{j}^{(i)}(B)$ (bottom) as illustrated in Figure 1. Note that $Q_{j}^{(i)}(M)=\emptyset$ if $i=1$. For any of these subsets, its cardinality is no more than $d$, and $|Q_{j}^{(i)}|\leq 3d$, i.e., every gate $U^{(i)}_{j}$ acts non-trivially on no more than $3d$ qubits.
@@
In summary, there are no more than $n/d+4$ rotations in total, since the rotations are in one-to-one correspondence with the two-layer reduced gates. Also notice that $\tilde{V}^{\prime 12\epsilon}_{j}$ and $\tilde{V}^{\prime 12\epsilon}_{k}$ overlap only if $|j-k|\leq 2$. Taking into account $\otimes_{j}\mathcal{P}_{j}^{28\epsilon}$, we conclude that the number of rotations acting nontrivially at a qubit is no more than 4. Finally, since a rotation on $8d$ qubits can be characterised by $14^{8d}-$ 1 parameters and there are $O(n/d)$ such rotations, the total number of real parameters of the local model is $O((n/d)\cdot 4^{8d})$.
IV Quantum local asymptotic normality (Q-LAN).
@@
In the following, we bridge the gap between the existing results and the setting of this work by finding a new Q-LAN for pure states generated by $O(\epsilon)$-rotations. For the generality of our results, let us consider a pure state generated by the concatenation of $G$ gates: $|\psi\rangle=U_{G}U_{G-1}\cdots U_{1}|0\rangle$, where each of the gates is an $\epsilon$-rotation that acts non-trivially only on a constant number of qubits. The reduced shallow-circuit state after parameter localization (see Section III) stands as a special case.
@@
Finally, directly applying Lemmas 5 and 6 yields that any $|\psi(\vec{H})\rangle\in\mathsf{S}_{N,\eta}$ is $O(\eta^{2}G^{2}/N)$-close (in trace distance) to
@@
where $h^{\vec{k}}:=\langle\vec{k}|\sum_{j=1}^{G}H_{j}|\vec{0}\rangle$. By Lemma 1, as long as $\eta^{2}G^{2}\ll\sqrt{N}$, the trace distance between $|\psi(\vec{H})\rangle\in\mathsf{S}_{N,\eta}$ and $|\psi^{\prime}(\vec{H})\rangle\in\mathsf{S}_{N,\eta}(Q)$ vanishes for large $N$. In this case, we can focus on constructing a Q-LAN for $|\psi^{\prime}(\vec{H})\rangle$.
@@
Noticing that $\sigma_{i}^{\vec{k}}\sigma_{j}^{\vec{k}^{\prime}}|0\rangle^{\otimes n}=0$ unless $\vec{k}=\vec{k}^{\prime}$, each locally-rotated state can be expanded as
@@
V A compression protocol for shallow-circuit states.
V.1 Preliminaries: truncation and amplification of coherent states.
As shown by the Q-LAN (Section IV), the $N$-copy states we are interested in are asymptotically equivalent to coherent states. Therefore, we introduce some preliminary protocols on processing coherent states, which will be used as building blocks of the compression protocol for shallow-circuit states.
@@
V.2 Efficient tomography of shallow-circuit states.
@@
V.3 A compression protocol for shallow-circuit states.
@@
Protocol 1 Compression protocol for $N$-copy shallow-circuit states.
@@
Our main result is that, when $\log_{2}N/\log_{2}n>12$, there exists a configuration $(\epsilon_{0},N_{0},\alpha_{0})$ such that Protocol 1 is a faithful compression protocol (i.e., one whose error vanishes as $N\to\infty$):
@@
Let $\mathsf{U}_{\rm sc}$ be the collection of all $n$-qubit pure states generated by depth-$d$ brickwork circuits, with $d$ being a fixed parameter. When $N=\Theta(n^{12+\gamma})$ for some $\gamma>0$, for any $\Delta\in\left(\frac{2}{12+\gamma},\frac{1}{2}-\frac{4}{12+\gamma}\right)$, there exist a compression protocol that faithfully compresses $N$ copies of any state from $\mathsf{U}_{\rm sc}$ into a hybrid memory of $8(1-\Delta)nd\left(\log_{2}N+O\left(\log_{2}n\right)\right)$ classical bits and $(n/d+4)2^{8d}\cdot\left(\Delta\log_{2}N+O(\log_{2}n)\right)$ qubits.
The proof is provided in Subsection V.4.
The total memory cost is in the order of $n\cdot\log_{2}N$. Specifically, the compression requires $8(1-\Delta)nd\log_{2}N+O\left(nd\log_{2}n\right)$ classical bits plus $(n/d+4)2^{8d}\cdot\left(\Delta\log_{2}N+O(\log_{2}n)\right)$ qubits. The quantum-to-classical cost ratio (in the leading order) is
$r_{\rm q-c}=\frac{2^{8d}\Delta}{8d^{2}(1-\Delta)}$ (79)
When the number of copies gets larger fixing $n$, i.e., when $\gamma\gg 1$, we may choose $\Delta\to 2/(12+\gamma)$ and the ratio $r_{\rm q-c}=O(1/\gamma)$, which means that the hybrid memory consists mainly of classical bits. In this sense, when $N$ is large enough, $N$-copy shallow-circuit states can be converted into classical bits plus a small portion of qubits. On the other hand, these qubits are indispensable as shown in Proposition 1. In this sense, the compressed $N$-copy state can be regarded as a long (classical) file with a quantum signature appended to it, similar as in the case of general qudits [33].
V.4 Proof of Theorem 3.
Here we prove the main theorem on the effectiveness of shallow-circuit state compression. We will argue that the following configuration $(\epsilon_{0},N_{0},\alpha_{0})$ of Protocol 1 achieves the desired features:
$\epsilon_{0}=n\cdot d\cdot N^{-\frac{1}{2}(1-\Delta)}\qquad N_{0}=N^{1-\frac{1}{2}\Delta}\qquad\alpha_{0}=896\sqrt{N}\epsilon_{0}.$ (80)
Faithfulness. By data processing and the triangle inequality, we can bound the overall error of the compression protocol by the sum of the errors of each step:
@@
    copies of $|\psi_{\rm sc}\rangle$. Since $\epsilon_{0}=nd/\sqrt{N^{1-\Delta}}$, we have $N^{\prime}_{0}=O\left(N^{1-\Delta}(\log N)^{2}\right)$ and thus $N_{0}=N^{1-\Delta/2}\gg N_{0}^{\prime}$, i.e., the number of copies allocated to tomography is sufficient to guarantee the desired accuracy. Notice that, when the estimate falls outside the confidence region, the distance between the estimate and the real state is upper bounded by one. Then the tomography error, $\epsilon_{\rm tomo}$, can be upper bounded by the probability that the estimate fails outside the confidence region:
@@
    $W_{j}^{112\epsilon_{0}}=\exp\left\{-i\frac{224ndN^{\Delta/2}}{\sqrt{N}}H_{j}\right\}$ (84)
    for some $(8d)$-partite Hamiltonian $H_{j}$ with $\|H_{j}\|_{\infty}\leq 1$. Since $224ndN^{\Delta/2}\cdot(n/d+4)\ll N^{\frac{1}{4}}$, the error of Q-LAN, $\epsilon_{\rm qlan}$, vanishes as guaranteed by Theorem 2.
  • The error of photon-number truncation (i.e., of compressing the multi-mode coherent state), $\epsilon_{\rm trun}$, can be bounded as follows. First, substituting $n_{\rm overlap}=4$ and $\eta=224ndN^{\Delta/2}$ into Eq. (44), the maximum amplitude of the multi-mode coherent state can be bounded as:
    $\max_{\vec{k}}|u^{\vec{k}}_{1}+iu^{\vec{k}}_{2}|\leq 896ndN^{\Delta/2}=896\sqrt{N}\epsilon_{0}.$ (85)
@@
    $\epsilon_{\rm trun}=O\left(n2^{8d}e^{-n^{2}d^{2}N^{\Delta}}\right),$ (86)
@@
  • The amplification error, $\epsilon_{\rm amp}$, can be bounded by directly substituting the intensity gain $N/(N-N_{0})$ into Lemma 7. The error for each single mode is thus $1-\sqrt{N-N_{0}/N}=O(N^{-\frac{1}{2}\Delta})$, and for the entire state we have

### v3

@@ Abstract
This work serves as the first step in establishing a new direction of coherent quantum information processing where the complexity of resources determines the rate and performance of processing, which goes beyond the existing literature that focused on incoherent information processing [9, 10, 11]. For future perspectives, it is our goal to consider more tasks such as cloning [19, 20, 21, 22] and gate programming [56, 57, 58, 59, 53] and, ultimately, to re-examine the entire quantum Shannon theory established in the past decade from the new perspective of the NISQ era.
Acknowledgement. The author acknowledges Penghui Yao and Fang Song for helpful discussions on pseudorandom states and Xinhui Yang for making Figure 1. This work is supported by the National Natural Science Foundation of China via the Excellent Young Scientists Fund (Hong Kong and Macau) Project 12322516, Guangdong Provincial Quantum Science Strategic Initiative (GDZX2303007 and GDZX2203001), Guangdong Basic and Applied Basic Research Foundation (Project No. 2022A1515010340), and the Hong Kong Research Grant Council (RGC) through the Early Career Scheme (ECS) grant 27310822 and the General Research Fund (GRF) grant 17303923.
End Matter on the error and the memory cost of Protocol 2. Here we provide an analysis of the total error and the memory cost of our compression protocol.
The total error will be vanishing if the error of each subroutine vanishes. First, the tomography, with an error that scales as $1/\sqrt{N_{0}}\ll\epsilon_{0}$, determines an estimate $\hat{U}|0\rangle^{\otimes n}$ such that $|\psi_{{\rm sc}}\rangle$ is within its $\epsilon_{0}$-neighborhood with high chance. Applying the inverse of $\hat{U}$ thereby localizes the remaining copies to a neighborhood with only $\mathsf{poly}(n)$ parameters. By Theorem 1, the generating circuit of the localized state consists of $O(n)$ $\epsilon_{0}$-rotations. Via the Q-LAN and the inverse Q-LAN (cf. Theorem 2), we can interchange between the remaining $N-N_{0}$ copies of the shallow-circuit state and a multi-mode coherent state $|\vec{u}\rangle_{\rm coh}$. The coherent state $|\vec{u}\rangle_{\rm coh}$ has $O(n)$ modes, whose amplitudes’ moduli are all upper bounded by $\eta=O(\sqrt{N}\epsilon_{0})$. By Theorem 2, the Q-LAN and the inverse Q-LAN have vanishing errors, as long as $\sqrt{N}\gg n^{2}N\epsilon_{0}^{2}$. The copies consumed by tomography are translated into the decrease of the coherent state’s amplitude by a ratio of $\sqrt{\frac{N}{N-N_{0}}}$, which can be compensated by amplitude amplification [60]. The idea here is similar as that of approximate and asymptotic cloning of quantum states. Although quantum states cannot be cloned perfectly [61, 62], they can be cloned approximately. In particular, for the task of generating $N_{0}$ extra copies out of $N-N_{0}$ input copies, the infidelity of cloning vanishes if $N\gg N_{0}\gg 1$ [20], which is the case here. Now, the coherent state has $O(n)$ modes. For each mode, we apply an individual amplifier. The total amplification error is ensured to be vanishing as long as the single amplifier’s error, which scales as $N_{0}/N$, is small compared to $1/n$. At last, each mode of the coherent state can be compressed by a photon number truncation. As the photon number of a single-mode coherent state $|u\rangle_{\rm coh}$ follows the Poisson distribution with mean $|u|^{2}$, it is enough to truncate at a suitable constant times the mean to ensure faithfulness, which can be satisfied by choosing $\alpha_{0}^{2}$ to be a suitable constant times $\epsilon_{0}^{2}N$.
As for the memory cost, it is stated in the protocol that the hybrid memory consists of two parts: a classical memory to store $\hat{U}$ and a quantum memory to store the truncated coherent state. The classical memory cost equals $\log_{2}|\mathsf{U}_{\rm sc}(\epsilon_{0})|$ bits, where $|\mathsf{U}_{\rm sc}(\epsilon_{0})|$ denotes the cardinality of the $\epsilon_{0}$-covering set. Intuitively, for a circuit of $G$ two-qubit gates, we can construct an $\epsilon_{0}$-covering by concatenating $G$ $(\epsilon_{0}/G)$-coverings of two-qubit gates. Each $\epsilon$-covering of two-qubit gates can be stored with $O(\log_{2}(1/\epsilon))$ bits, and thus the cost for the $G$ gate covering scales as $G\log_{2}(G/\epsilon_{0})$. In our case, $G=O(n)$ for shallow circuits and, since $\epsilon_{0}=\Theta(n\cdot N^{-\frac{1}{2}(1-\frac{2}{3}\Delta)})$, the size of the classical memory is $O(n(1-\frac{2}{3}\Delta)\log_{2}N)$. For the quantum memory, each mode of the coherent state has an amplitude upper bounded by $O(\sqrt{N}\epsilon_{0})$, according to Eq. (3) and Theorem 1. Since there are $O(n)$ modes and each mode’s photon number is truncated at $O(\alpha_{0}^{2})=O(n^{2}N^{\Delta})$, it requires $O(\Delta\log_{2}N)$ qubits (assuming $N\gg n$) to store each mode, and $O(n\Delta\log_{2}N)$ qubits in total.
We note that the aforementioned constraints ($\sqrt{N}\gg n^{2}N\epsilon_{0}^{2}$ and $N_{0}/N\ll 1/n$) together set a constraint on the relation between $n$ and $N$, as state in Theorem 6.
Appendix A Preliminaries
A.1 Conventions and notations
@@
A.2 Distance measures for quantum states and gates
@@
where $\mathcal{H}_{A}\simeq\mathcal{H}$ is an ancillary space. For two unitary channels $\mathcal{U}$ and $\mathcal{V}$, the diamond norm is well-captured by the operator norm (cf. [63, Proposition 1.6]):
@@
A.3 Properties of $N$-copy states.
@@
Therefore, the task of $N$-copy state compression cannot be trivially reduced to quantum state tomography. In Appendix D.6, we will show the same result for shallow-circuit states.
Appendix B Efficient local parameterization of shallow-circuit states
In Section D we will see that, to achieve faithful compression, we need to pin the shallow-circuit state $|\psi_{{\rm sc}}\rangle$ down to a small local region parameterised by up to $\mathsf{poly}(n)$ real parameters. We name such a task as local parameterization of quantum states, which is achieved by measuring copies of $|\psi_{{\rm sc}}\rangle$ and then constructing the local model. However, it is more challenging than the task of quantum state tomography, as one has to not only identify a neighborhood of the state but also parameterise this neighborhood efficiently.
B.1 Local parameterization of quantum states
@@
Lemma 2 (Local parameterization of $\epsilon$-rotations).
@@
Definition 3 (Local parameterization of quantum states).
Given a set $\mathsf{S}=\{|\psi\rangle\}$ of $n$-qubit quantum states and a fixed quantum state $|\psi_{0}\rangle=U_{0}|0\rangle^{\otimes n}$ with the guarantee that $d_{\operatorname{Tr}}(\psi,\psi_{0})\leq\epsilon$ for every $|\psi\rangle\in\mathsf{S}$, the task of local parameterization of $\mathsf{S}$ is to identify a set $\mathsf{W}_{\rm loc}$ of $O(\epsilon)$-rotations, such that for every $|\psi\rangle\in\mathsf{S}$ there exists a $W\in\mathsf{W}_{\rm loc}$ with $|\psi\rangle=U_{0}W|0\rangle^{\otimes n}$. Moreover, the local parameterization is efficient if $\mathsf{W}_{\rm loc}$ can be parameterized with $\mathsf{poly}(n)$ local parameters, i.e., if there exists an on-to function $\mathcal{P}:\mathsf{E}^{f}\to\mathsf{W}_{\rm loc}$ with $\mathsf{E}\subset\mathbb{R}$ being an $O(\epsilon)-$ length interval and $f\in\mathsf{poly}(n)$.
@@
The above “vanilla” local parameterization of quantum states appears a simple task. However, making it parameter-efficient is not a trivial task. The local parameterization in Lemma 3 consists of generic $n$-qubit unitaries and has exponentially many degrees of freedom.
In practice, it is more favorable to seek an efficient localization. A hope is when the set of states under consideration has certain structure. For instance, when $\mathsf{S}$ consists of states generated by acting a low-dimensional unknown quantum gate upon a fixed state. If all such states in $\mathsf{S}$ are close to the original fixed state, it is quite natural to speculate that we could probably replace this unknown unitary by an $O(\epsilon)$-rotation to achieve an efficient local parameterization (since the $O(\epsilon)$-rotation is also low-dimensional). Somewhat surprisingly, this intuition turns out to be wrong:
@@
In general, it is not possible to partially locally parameterize a quantum state. That is, there exists a bipartite pure state $|\psi\rangle_{AB}$ and a unitary $\mathcal{V}_{B}$ acting on $B$ such that:
@@
The purpose of the above remark is to justify that our goal – an efficient local parameterization of shallow-circuit states – is not as simple as it appears. One of our main results is to rigorously prove that such an efficient local parameterization indeed exists for shallow-circuit states.
Theorem 4 (Efficient local parameterization of shallow-circuit states.).
@@
Here $W=\prod_{j=1}^{n_{\rm gate}}W^{28\epsilon}_{j}$ with $n_{\rm gate}\leq n/d+4$, and each $W^{28\epsilon}_{j}$ is a $(28\epsilon)$-rotation acting on no more than $8d$ qubits. Among $\{W_{j}^{28\epsilon}\}$, the maximum of the number of gates that act non-trivially on the same qubit is upper bounded by $n_{\rm overlap}=4$. The set $\mathsf{W}_{\rm loc}$ can be parameterized with $O(n\cdot(2^{16d}/d))$ (real) local parameters.
Note that the bound on the number of parameters can be made tighter (but the bound might become less succinct). Details can be found in the proof; see Appendix B.2.
@@ B.2 Proof of Theorem 4
Grouping qubits. For convenience of discussion, we group the $n$ qubits into subsets and label them using the following convention. We define a function $Q=Q(U)$ that maps a unitary $U$ to the qubits that it acts non-trivially on. In particular, $Q_{j}^{(i)}:=Q(U_{j}^{(i)})$ is the collection of qubits that $U_{j}^{(i)}$ acts non-trivially on. We further divide $Q_{j}^{(i)}$ into subsets $Q_{j}^{(i)}(T)$ (top), $Q_{j}^{(i)}(M)$ (middle) and $Q_{j}^{(i)}(B)$ (bottom) as illustrated in Figure 2. Note that $Q_{j}^{(i)}(M)=\emptyset$ if $i=1$. For any of these subsets, its cardinality is no more than $d$, and $|Q_{j}^{(i)}|\leq 3d$, i.e., every gate $U^{(i)}_{j}$ acts non-trivially on no more than $3d$ qubits.
@@
In summary, there are no more than $n/d+4$ rotations in total, since the rotations are in one-to-one correspondence with the two-layer reduced gates. Also notice that $\tilde{V}^{\prime 12\epsilon}_{j}$ and $\tilde{V}^{\prime 12\epsilon}_{k}$ overlap only if $|j-k|\leq 2$. Taking into account $\otimes_{j}\mathcal{P}_{j}^{28\epsilon}$, we conclude that the number of rotations acting non-trivially at a qubit is no more than 4. Finally, since a rotation on $8d$ qubits can be characterized by $14^{8d}-$ 1 parameters and there are $O(n/d)$ such rotations, the total number of real parameters of the local model is $O((n/d)\cdot 4^{8d})$.
Appendix C Quantum local asymptotic normality (Q-LAN).
@@
In the following, we bridge the gap between the existing results and the setting of this work by finding a new Q-LAN for pure states generated by $O(\epsilon)$-rotations. For the generality of our results, let us consider a pure state generated by the concatenation of $G$ gates: $|\psi\rangle=U_{G}U_{G-1}\cdots U_{1}|0\rangle$, where each of the gates is an $\epsilon$-rotation that acts non-trivially only on a constant number of qubits. The reduced shallow-circuit state after parameter localization (see Appendix B) stands as a special case.
@@
Finally, directly applying Lemmas 5 and 6 yields that any $|\psi(\vec{H})\rangle$ is $O(\eta^{2}G^{2}/N)$-close (in trace distance) to
@@
where $h^{\vec{k}}:=\langle\vec{k}|\sum_{j=1}^{G}H_{j}|\vec{0}\rangle$. By Lemma 1, as long as $\eta^{2}G^{2}\ll\sqrt{N}$, the trace distance between $|\psi(\vec{H})\rangle^{\otimes N}\in\mathsf{S}_{N,\eta}(Q)$ and $|\psi^{\prime}(\vec{H})\rangle^{\otimes N}$ vanishes for large $N$. In this case, we can focus on constructing a Q-LAN for $|\psi^{\prime}(\vec{H})\rangle$.
@@
Noticing that $\sigma_{i}^{\vec{k}}\sigma_{j}^{\vec{k}^{\prime}}|0\rangle^{\otimes n}=0$ unless $\vec{k}=\vec{k}^{\prime}$, each locally rotated state can be expanded as
@@
Appendix D A compression protocol for shallow-circuit states.
D.1 Preliminaries: truncation and amplification of coherent states.
As shown by the Q-LAN (Appendix C), the $N$-copy states we are interested in are asymptotically equivalent to coherent states. Therefore, we introduce some preliminary protocols on processing coherent states, which will be used as building blocks of the compression protocol for shallow-circuit states.
@@
D.2 Efficient tomography of shallow-circuit states.
@@
D.3 A compression protocol for shallow-circuit states.
@@
Protocol 2 Compression protocol for $N$-copy shallow-circuit states.
@@
Our main result is that, when $\log_{2}N/\log_{2}n>32/3$, there exists a configuration $(\epsilon_{0},N_{0},\alpha_{0})$ such that Protocol 2 is a faithful compression protocol (i.e., one whose error vanishes as $N\to\infty$):
@@
Let $\mathsf{U}_{\rm sc}$ be the collection of all $n$-qubit pure states generated by depth-$d$ brickwork circuits, with $d$ being a fixed parameter. When $N=\Theta(n^{32/3+\gamma})$ for some $\gamma>0$, for any $\Delta\in\left(\frac{6}{32+3\gamma},\frac{3}{4}-\frac{18}{32+3\gamma}\right)$, there exist a compression protocol that faithfully compresses $N$ copies of any state from $\mathsf{U}_{\rm sc}$ into a hybrid memory of $8(1-2\Delta/3)nd\left(\log_{2}N+O\left(\log_{2}n\right)\right)$ classical bits and $(n/d+4)2^{8d+1}/3\cdot\left(\Delta\log_{2}N+O(\log_{2}n)\right)$ qubits.
The proof is provided in Appendix D.4.
The total memory cost is in the order of $n\cdot\log_{2}N$. The quantum-to-classical cost ratio (in the leading order) is
$r_{\rm q-c}=\frac{2^{8d-2}\Delta}{d^{2}(3-2\Delta)}$ (79)
When the number of copies gets larger fixing $n$, i.e., when $\gamma\gg 1$, we may choose $\Delta\to 6/(32+3\gamma)$ and the ratio $r_{\rm q-c}=O(1/\gamma)$, which means that the hybrid memory consists mainly of classical bits.
D.4 Proof of Theorem 6.
Here we prove the main theorem on the effectiveness of shallow-circuit state compression. We will argue that the following configuration $(\epsilon_{0},N_{0},\alpha_{0})$ of Protocol 2 achieves the desired features:
$\epsilon_{0}=n\cdot d\cdot N^{-\frac{1}{2}(1-\frac{2}{3}\Delta)}\qquad N_{0}=N^{1-\frac{1}{2}\Delta}\qquad\alpha_{0}=896\sqrt{N}\epsilon_{0}.$ (80)
Faithfulness. By data processing and the triangle inequality, we can upper-bound the overall error of the compression protocol by the sum of the errors of each step:
@@
    copies of $|\psi_{\rm sc}\rangle$. Since $\epsilon_{0}=nd/\sqrt{N^{1-2\Delta/3}}$, we have $N^{\prime}_{0}=O\left(N^{1-2\Delta/3}(\log N)^{2}\right)$ and thus $N_{0}=N^{1-\Delta/2}\gg N_{0}^{\prime}$, i.e., the number of copies allocated to tomography is sufficient to guarantee the desired accuracy. Notice that, when the estimate falls outside the confidence region, the distance between the estimate and the real state is upper bounded by one. Then the tomography error, $\epsilon_{\rm tomo}$, can be upper bounded by the probability that the estimate fails outside the confidence region:
@@
    $W_{j}^{112\epsilon_{0}}=\exp\left\{-i\frac{224ndN^{\Delta/3}}{\sqrt{N}}H_{j}\right\}$ (84)
    for some $(8d)$-partite Hamiltonian $H_{j}$ with $\|H_{j}\|_{\infty}\leq 1$. Since $224ndN^{\Delta/3}\cdot(n/d+4)\ll N^{\frac{1}{4}}$, the error of Q-LAN, $\epsilon_{\rm qlan}$, vanishes as guaranteed by Theorem 5.
  • The error of photon-number truncation (i.e., of compressing the multi-mode coherent state), $\epsilon_{\rm trun}$, can be bounded as follows. First, substituting $n_{\rm overlap}=4$ and $\eta=224ndN^{\Delta/3}$ into Eq. (44), the maximum amplitude of the multi-mode coherent state can be bounded as:
    $\max_{\vec{k}}|u^{\vec{k}}_{1}+iu^{\vec{k}}_{2}|\leq 896ndN^{\Delta/3}=896\sqrt{N}\epsilon_{0}.$ (85)
@@
    $\epsilon_{\rm trun}=O\left(n2^{8d}e^{-n^{2}d^{2}N^{2\Delta/3}}\right),$ (86)
@@
  • The amplification error, $\epsilon_{\rm amp}$, can be bounded by directly substituting the intensity gain $N/(N-N_{0})$ into Lemma 7. The error for each single mode is thus $N_{0}/N=O(N^{-\frac{1}{2}\Delta})$, and for the entire state we have